OpenStudy (anonymous):

If g(x ) = 3x + 1, evaluate g(-14). -41 -10 -5

OpenStudy (anonymous):

@joyraheb

OpenStudy (anonymous):

g(-14) = 3(-14) + 1= -42+1 = -41